Explain in detail, with the aid of a graph, the two components of the demand for money. Identify the main determinant of each component. In your graph, illustrate each component as well as the total demand for money.
1
Expert's answer
2020-09-02T11:30:52-0400

Demand for money is the total amount of money that a population holds in terms of financial assets. Two components are discussed below for holding money.

i) Precautionary purpose

ii) Transaction purpose

  • Precautionary Motive, is the tendency of a firm or people to hold cash for emergencies or new business opportunities i.e sickness,fires etc these occurrences need cash in order for them to be cushioned.This type of demand for money is determined by income and the general level of business activity. Precautionary demand is income-elastic, it is expressed as "M^P=f(Y)" where "M^P" is the precautionary demand for money and "f(Y)" is the function of income.
  • Transaction purpose, it entails holding money in to carry out day to day functions which can be commercial in nature.Transaction motive for money is also determined by the level of income. This is depicted by the equation below;
"M^T=kY\\
Where\ M^T=Transaction\ demand\ for\ money\\
Y=National\ income\\
k=constant\ of\ proportionality,\ 0<k<1"

Then we will combine the transaction and precautionary motive into a single demand schedule which is known as demand for active balances, which represents the total demand for money. Let "M_A" represent active balances, we can therefore say;


"M_A=M^P+M^T\\
Substitute\ eq.(1) into\ eq.\ (2), we\ then\ have\\
M_A=M^p+kY"

A plot of equation. (3) is a straight line; that is, the "M_A" demand for money is equal to constant "M^P" demand for money is equal to a constant "M^P" plus a coefficient "k" times a variable "Y" and such an equation shows a straight line,as in the graph below.